CARNEGIE MELLON UNIVERSITY DEPARTMENT OF MATHEMATICAL SCIENCES

CENTER FOR NONLINEAR ANALYSIS

SEMINAR SCHEDULE FOR ACADEMIC YEAR 2001/2002

TUESDAY, September 25, 2001

CNA SEMINAR: 1:30 P.M., PPB 300, Andreas Veeser, Università degli Studi di Milano. TITLE:

MONDAY, July 23, 2001

CNA SEMINAR: 1:30 P.M., PPB 300, H. Christian Gromoll, University of California, San Diego. TITLE: “Diffusion Approximation for a Processor Sharing Queue in Heavy Traffic” ABSTRACT: We discuss a diffusion approximation result for the queue length process of a GI/GI/1 processor sharing queue in heavy traffic. The main approach is to study a certain measure valued process, which encodes information about the residual service times and from which standard measures of performance such as queue length, workload and sojourn time can be recovered.